PDiddi:

Hey! I am a sophomore and I am doing geometry. This question is about circles and their equation. as you might know...I guess...their equation is (x−h)^2+(y−k)^2=r^2 So this is my problem, it says to WRITE the equation of the circle with the given information: Center: (6,4) Tangent to y = -8 Can anyone please help?

darkknight:

Figure out what the h, k and r is when the Line is tangent to the curve. H and K is given

PDiddi:

Yes, I can see that, but I do not understand where I can input the Tangent to y.

surjithayer:

correction difference between (6,4) and (6,-8) \[r=\sqrt{(6-6)^2+(-8-4)^2}=\sqrt{(-12)^2}=12\] eq. of circle is \[(x-6)^2+(y-4)^2=12^2\] \[(x-6)^2+(y-4)^2=144\]